OpenStudy (anonymous):

Architect Frank Lloyd Wright included a pool shaped like a right triangle in his design of Taliesin West in Scottsdale, Arizona. The sides of the pool that represent the legs measure 41 ft and 40 ft. Calculate the perimeter of the pool. Round the answer to the nearest foot.

OpenStudy (anonymous):

To find the other side of the triangle: (Other)^2 = (41^2) + (40^2) Other = ? Perimeter of pool = Other + 41 + 40